OpenStudy (anonymous):

Factor when possible: -x^2y+3xy+70y= (type your answer in Factored Form. i am not sure how to do this, any help would be appriciated.

OpenStudy (anonymous):

-y(x^2-3x-70) check if this is the answer

OpenStudy (anonymous):

i have no idea how to check if that is the answer.

OpenStudy (anonymous):

:) ok never mind because i checked it before i posted it, what i wanted you to do (and this is only if this is from a textbook) was to look in the back and see if that is the answer that's there

OpenStudy (anonymous):

and to check it you would just multiply it

OpenStudy (anonymous):

-x^2y+3xy+70y=-y(x^2-3x-70)=-y(x=10)(x+7)

OpenStudy (anonymous):

how did you get that last part?

OpenStudy (anonymous):

Sorry messed up on an equals sign. -x^2y+3xy+70y=-y(x^2-3x-70)=-y(x-10)(x+7). You know it's (x-10) and (x+7) because when we look at (x^2-3x-70), we want two numbers that multiply to give -70 (the constant term) and add to give -3 (the coefficient on the x). These numbers are -10 and 7. So (x^2-3x-70) factors as (x-10)*(x+7)

OpenStudy (anonymous):

oh yeah, i missed that part